Problem on genetics, Please solve for what is asking for the first and the second pedigree, I have included every single detail that is helpful to the problem

Antoine’s narrative:

I have a sister who has sickle cell disease. She is not married. Neither my mom nor my dad has the disease. My mother has two brothers and one sister, although one of her brothers passed away when I was very young from complications of sickle cell anemia. Neither my uncle nor my aunt has the disease; they are both married to unaffected spouses, and each has a daughter. None of my cousins are affected. My maternal grandparents were both normal; they were from relatively large families (grandmother had five siblings, grandfather had six), but no one in the family seems to know much about the health status of my great uncles and aunts. At the very least, there’s no immediate indication that any of them suffered from sickle cell.

My father is normal. He has a brother who is also normal, who is married to an unaffected wife. They have three children (two daughters and a son), none of whom are affected. My paternal grandparents were both normal, but my grandmother had a brother who died of sickle cell anemia. She also had a sister who was unaffected. My grandfather had four siblings, none of whom were affected. My parents are still alive, but only my paternal grandmother is still alive.

Taisha’s narrative:

I have a sister who does not display symptoms of sickle cell. She is married to an unaffected man, but they haven’t yet had children. Neither of my parents is affected. However, my mom’s younger sister died about ten years ago from sickle cell anemia. She did get married, though, to an unaffected man, and has a son who is normal. My mother also has a brother who seems normal, but who once suffered sickle cell like symptoms when he was hiking Mount Kilimanjaro. A friend who knows about this suggested that this meant he was a carrier of the disease. He has a normal wife, and a son and a daughter, both of whom are unaffected. My maternal grandfather was normal, but my maternal grandmother had sickle cell. She was really tough, though, and lived into her fifties. She had two brothers and two sisters, none of whom had sickle cell; I wasn’t able to find out about my paternal grandfather’s siblings. My father is normal, but his father had sickle cell. He died soon after my father was born. My paternal grandmother is normal, and is still alive. She remarried several years after my grandfather died.

Genetic counsellor’s pedigree:

Problem on genetics, Please solve for what is aski

Using the information provided in the narratives, along with the pedigree above, deduce the pattern of inheritance for sickle cell anemia. Employing the shorthand ‘A’ for the HbA allele and ‘S’ for the HbS allele, assign as many genotypes as you can to the individuals in the pedigree above.

Using the information you have amassed about Antoine and Taisha’s family histories, estimate the probability that they would have a child that suffers from full sickle cell anemia.

Probability: ________________________________________

Optional Challenge Problem:

Imagine that Antoine and Taisha gave the following narratives, instead of those provided above. Using these, and the other pieces of information provided, answer the following questions.

Antoine’s narrative:

I have a sister who has sickle cell disease. She is not married. Neither my mom nor my dad has the disease. My mother has two brothers and one sister, although one of her brothers passed away when I was very young from complications of sickle cell anemia. Neither my uncle nor my aunt has the disease; they are both married to unaffected spouses, and each has a daughter. None of my cousins are affected. My maternal grandparents were both normal; they were from relatively large families (grandmother had five siblings, grandfather had six), but no one in the family seems to know much about the health status of my great uncles and aunts. At the very least, there’s no immediate indication that any of them suffered from sickle cell.

My father is normal. He has a brother who is also normal, who is married to an unaffected wife. They have three children (two daughters and a son), none of whom are affected. My paternal grandparents were both normal, but my grandmother had a brother who died of sickle cell anemia. She also had a sister who was unaffected. My grandfather had four siblings, none of whom were affected. My parents are still alive, but only my paternal grandmother is still alive.

Taisha’s narrative:

I have a sister who does not display symptoms of sickle cell. She is married to an unaffected man, but they haven’t yet had children. Neither of my parents is affected. However, my mom’s younger sister died about ten years ago from sickle cell anemia. She did get married, though, to an unaffected man, and has a son who is normal. My mother also has a brother who seems normal, but who once suffered sickle cell like symptoms when he was hiking Mount Kilimanjaro. A friend who knows about this suggested that this meant he was a carrier of the disease. He has a normal wife, and a son and a daughter, both of whom are unaffected. My maternal grandparents were both unaffected; my grandmother had two brothers and two sisters, none of whom had sickle cell, and my grandfather had three brothers and two sisters, all of whom were also normal. My father is normal, but he was adopted from an orphanage, and nothing is known about his family.

Genetic counsellor’s pedigree:

Using the information provided in the narratives, along with the pedigree above, deduce the pattern of inheritance for sickle cell anemia at the organismal level (the obvious outward phenotype), at the cellular level, and at the molecular level, and provide your rationale for making these decisions.

Hint: The organismal phenotype should be relatively easy to deduce, but, to determine the cellular and molecular phenotypes, you should give special consideration to Taisha’s uncle, along with what is known about him from the narrative. What could possibly have led to him developing sickle cell like symptoms while hiking?
Answer #1

Case 1: The following is the pedigree for the given information:

AS AS AS ASAS AASS AS AS AS AS AS AS Taisha ntonie

The probability of Taisha and Antony's child suffer from sickle cells is calculated as given below:

The probability that both have the genotypes SS x SS = 1/4 x 1/4 = 1/16 = 0.0625

The probability that both are heterozygous is AS x AS = 1/2 x 1/2 = 1/4. If both are heterozgygous, 1/4 of the progeny probably inherit the condition. Therefore, the probability is 1/4 x 1/4 = 0.0625

In no other combinations does there occurs a change that the child is affected. Therefore, the probability that the child could be affected is 0.0625 + 0.0625 = 0.125